OpenStudy (anonymous):

Solve for x: 2 |x - 3| + 4 = 16

OpenStudy (anonymous):

\[2|x-3|+4=16\] \[2|x-3|=12\] \[|x-3|=6\] \[x-3=6\] or \[x-3=-6\] \[x=9\] or \[x=-3\]